# Satisficing Regret Minimization in Bandits

### v2

@@
Satisficing Regret Minimization in Bandits
@@
Motivated by the concept of satisficing in decision-making, we consider the problem of satisficing regret minimization in bandit optimization. In this setting, the learner aims at selecting satisficing arms (arms with mean reward exceeding a certain threshold value) as frequently as possible. The performance is measured by satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. We propose SELECT, a general algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence bound Testing, that attains constant satisficing regret for a wide variety of bandit optimization problems in the realizable case (i.e., a satisficing arm exists). Specifically, given a class of bandit optimization problems and a corresponding learning oracle with sub-linear (standard) regret upper bound, SELECT iteratively makes use of the oracle to identify a potential satisficing arm with low regret. Then, it collects data samples from this arm, and continuously compares the LCB of the identified arm’s mean reward against the threshold value to determine if it is a satisficing arm. As a complement, SELECT also enjoys the same (standard) regret guarantee as the oracle in the non-realizable case. Finally, we conduct numerical experiments to validate the performance of SELECT for several popular bandit optimization settings.
@@
online learning, bandits, satisficing, regret
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic framework for decision-making under uncertainty. In MAB, a learner is given a set of arms with initially unknown mean rewards. She repeatedly picks an arm to collect its noisy reward and her goal is to maximize the cumulative rewards. The performance is usually measured by the notion of regret, which is the difference between the maximum possible total mean rewards and the total mean rewards collected by the learner. With its elegant form, MAB has found numerous applications in recommendation systems, ads display, and beyond, and many (nearly) optimal algorithms have been developed for various bandit optimization settings (Lattimore and Szepesvári 2020).
@@
The concept of satisficing finds its place in many real world decision-making under uncertainty settings (Caplin et al. 2011). For example, in inventory management, studies show that product stockouts can result in customer disengagement and impair a firm’s profitability (Fitzsimons 2000, Jing and Lewis 2011). As such, retailers usually replenish their products periodically to improve service level (i.e., probability of no stockout for any product). Notably, due to various constraints (e.g., capacity and inventory cost), instead of blindly maximizing service level, retailers typically set a target service level (e.g., the ideal service level in retailing is $90\%\sim 95\%$ (Levy et al. 2023)). Consequently, in face of the unknown demand distributions, one can cast the problem into a MAB setting with a different objective. In this setting, a retailer starts without complete knowledge of the demand distributions and proceeds by trying various stock level combinations (i.e., arms) of the products over time. Here, the objective is to hit the target service level as frequently as possible rather than maximizing the service level every time. As another example, in fashion product design, designers usually need to design a sequence of products for a group of customers with initially unknown preferences. Instead of demanding for the best design every time, designers only need to pick a design that can hopefully meet the expectation of a certain portion of the customers for most of the time (Mabey 2022, Whitenton 2014).
@@
The above examples give rise to the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2023) where the objective is to hit a satisficing arm (i.e., an arm whose mean reward exceeds a certain threshold value) as frequently as possible. The corresponding performance measure is satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. In particular, Michel et al. (2023) develops SAT-UCB, an algorithm for finite-armed satisficing bandits: In the realizable case (i.e., a satisficing arm exists), the algorithm attains a constant satisficing regret, i.e., independent of the length of the time horizon; Otherwise, in the non-realizable case (i.e., the threshold value is out of reach), it preserves the optimal logarithmic (standard) regret.
@@
Despite these, existing algorithms usually impose that there is a clear separation between the threshold value and the largest mean reward among all non-satisficing arms, which is referred to as the satisficing gap. Moreover, the satisficing regret bounds usually scale inverse proportionally to the satisficing gap. On the other hand, it is evident that in many bandit optimization problems (e.g., Lipschitz bandits and concave bandits), the satisficing gap can simply be 0, making the bounds vacuous. Therefore, it is not immediately clear if SAT-UCB (Michel et al. 2023) or other prior results can go beyond finite-armed bandits and if constant satisficing regret is possible for bandit optimization in general.
@@
Main Contributions: In this work, we propose SELECT, a novel algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence Bound Testing. More specifically:
@@
  • We describe the design of SELECT in Section 3. For a given bandit optimization problem class and an associated learning oracle with sub-linear (but not necessarily optimal) standard regret guarantee, SELECT runs in rounds with the help of the oracle. At the beginning of a round, SELECT first applies the oracle for a number of time steps and randomly samples an arm from its trajectory as a candidate satisficing arm. Then, it conducts forced sampling by pulling this arm for a pre-specified number of times. Finally, it continues to pull this arm and starts to monitor the resulted lower confidence bound (LCB) of its mean reward. SELECT terminates the current round and starts the next once the LCB falls below the threshold value;
@@
  • In Section 4, we establish that in the realizable case or when a satisficing arm exists, SELECT is able to achieve a constant satisficing regret. Notably, the satisficing regret bound has no dependence on the satisficing gap (see the forthcoming Remark 2 for a detailed discussion). Instead, it scales inverse proportionally to the exceeding gap, which measures the difference between the optimal reward and the threshold value, and is positive in general; Otherwise (i.e., when no satisficing arm exists), SELECT enjoys the same standard regret bound as the learning oracle.
@@
  • Finally, in Section 6 we conduct numerical experiments to demonstrate the performance of SELECT in fnite-armed bandits, concave bandits, and Lipschitz bandits. We use SAT-UCB, SAT-UCB+ (a heuristic with no regret guarantee proposed by Michel et al. (2023)), and the respective learning oracle as the benchmarks. Our results reveal that in the realizable case, SELECT does achieve constant satisficing regret. Moreover, both its satisficing regrets (in the realizable cases) and standard regrets (in the non-realizable cases) either significantly outperform the benchmarks’ or are close to the best-performing ones’.
@@ 2 Problem Formulation
In this section, we present the setup of our problem. Consider a class of bandit optimization problem $(\mathcal{X},\mathcal{R})$, where $\mathcal{X}$ is the arm set and $\mathcal{R}\subseteq\{r:\mathcal{X}\to\mathbb{R}\}$ is the class of admissible reward functions. The learner is given a set of feasible arms $\mathcal{X}$ and a class of admissible reward functions $\mathcal{R}$, pulling arm $X\in\mathcal{X}$ in a time step brings a random reward with mean $r(X)$. The learner knows the underlying reward function $r(\cdot)$ belongs to the function class $\mathcal{R}$, but the exact $r(\cdot)$ is initially unknown to the learner.
In each time step $t=1,2,\ldots,T$, the learner pulls an arm $X_{t}\in\mathcal{X}$, and then observes and receives a noisy reward $Y_{t}=r(X_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is a conditionally 1-subgaussian random variable, i.e., $\mathbb{E}[\epsilon_{t}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]=0$ and $\mathbb{E}[e^{\lambda\epsilon_{t}}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]\leq e^{\lambda^{2}/2}$ holds for all $\lambda\in\mathbb{R}$.
@@
$\texttt{Regret}_{S}=\mathbb{E}\left[\sum_{t=1}^{T}\max\{S-r(X_{t}),0\}\right].$
Because it is possible that no arm achieves the satisficing level, we also follow the conventional bandit literature to define the (standard) regret, which measures the expected cumulative difference of the mean rewards between the optimal arm and the chosen arm’s, i.e., let $X^{*}=\arg\max_{X\in\mathcal{X}}r(X),$ the regret is defined as
$\texttt{Regret}=T\cdot r(X^{*})-\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t})\right].$
We say that the setting is realizable if $r(X^{*})\geq S$; Otherwise, it is non-realizable. We recall that an arm $X$ is satisficing if $r(X)\geq S$; Otherwise, it is non-satisficing. Existing works on similar settings (e.g., Michel et al. 2023) usually derive satisficing regret bound that scales inversely proportional to the satisficing gap
$\Delta_{S}=\min\{S-r(X):r(X)<S\}.$
@@ 3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
In this section, we present our algorithmic template SELECT for general bandit problems.
When the arm space $\mathcal{X}$ is not too large, one can collect enough data samples for every arm to estimate its mean reward and gradually abandon any non-satisficing arms (see, e.g., Garivier et al. 2019, Michel et al. 2023). In general, however, the arm space can be large and may even contain infinitely many arms (e.g., continuous arm space). Therefore, it becomes impossible to identify all non-satisficing arms. We thus take an alternative approach: It repeatedly locates a potentially satisficing arm with low (satisficing) regret. Then, it tests if this candidate arm is truly a satisficing arm or not. If not, it kicks off the next round of search.
To ensure fast detection of candidate arms with low regret for a given bandit problem class $(\mathcal{X},\mathcal{R})$, SELECT makes use of a bandit algorithm for (standard) regret minimization on $(\mathcal{X},\mathcal{R})$. To this end, we assume access to a blackbox learning oracle that achieves sub-linear standard regret for the bandit problem class $(\mathcal{X},\mathcal{R})$.
@@
For problem class $(\mathcal{X},\mathcal{R})$, there exists a sequence of learning algorithms ALG such that for any reward function $r\in\mathcal{R}$ and any time horizon $t\geq 2$, the regret of $\texttt{ALG}(t)$ is bounded by
$\texttt{Regret}=\mathbb{E}\left[t\cdot r(X^{*})-\sum_{s=1}^{t}r(X_{s})\right]\leq C_{1}t^{\alpha}\log(t)^{\beta}.$
@@
We remark that Condition 1 only asks for sub-linearity in standard regret upper bounds rather than optimality. Therefore, it is an extremely mild condition and can be easily satisfied by a wide range of bandit optimization problems and the corresponding bandit learning oracles, e.g., finite-armed bandits, Lipschitz bandits, etc. We will demonstrate this in the forthcoming Section 5 and Appendix D.
With this, for a problem class $(\mathcal{X},\mathcal{R})$ and its sub-linear regret learning algorithm ALG, SELECT runs in rounds. In round $i$, it takes the following steps (see Fig. 1 for a illustration):
@@
Step 1. Identify a Potentially Satisficing Arm: Let $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$, SELECT follows $\texttt{ALG}(t_{i})$ for the first $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil=\lceil 2^{i/\alpha}\rceil$ steps of round $i$ and records its arm selections. Then, it samples an arm $\hat{X}_{i}$ uniformly random from the trajectory of $\texttt{ALG}(t_{i})$ in this round. By virtue of ALG, we find an arm whose mean reward is at most $\tilde{O}(t^{\alpha-1}_{i})=\tilde{O}(\gamma_{i})$ below $r(X^{*})$ in expectation, and only $\tilde{O}(t^{\alpha}_{i})$ satisficing regret is incurred in this step in the realizable case. We note that in the realizable case, as $i$ increases, $\tilde{O}(\gamma_{i})$ will gradually become smaller than $\Delta_{S}^{*}$, meaning that the sampled arm $\hat{X}_{i}$ is more likely to be a satisficing arm and enjoys no satisficing regret;
Step 2. Forced Sampling on the Identified Arm: To validate if $\hat{X}_{i}$ is a satisficing arm or not, we need to collect enough data from pulling $\hat{X}_{i}$ before performing any statistical test (see the forthcoming Remark 2 for the reasons). However, when the arm set $\mathcal{X}$ is large, $\texttt{ALG}(t_{i})$ might only have pulled $\hat{X}_{i}$ for limited number of times. To circumvent this, we perform a forced sampling on $\hat{X}_{i}$ by pulling it for $T_{i}=\lceil\gamma^{-2}_{i}\rceil$ times;
Step 3. Lower Confidence Bound Tests: In this step, SELECT kicks off the LCB test as more data is collected from pulling $\hat{X}_{i}$. We use $k$ (initialized to 0) to denote the number of additional pieces of noisy reward data generated by pulling $\hat{X}_{i}$ in the current time step and $\hat{r}^{\text{tot}}_{i}$ to denote the running total reward collected from Step 2 and the current step so far. At the same time, it keeps comparing the LCB of $\hat{X}_{i}$’s mean reward, i.e., $\hat{r}^{\text{tot}}_{i}/(T_{i}+k)-\sqrt{4\log(T_{i}+k)/(T_{i}+k)}$, against the threshold value $S$ after acquiring every piece of new data. SELECT terminates this round and enters the next whenever the LCB is less than $S$.
@@
Remark 2.
(Key Novelties of SELECT) Our algorithm brings together several key ingredients, including sampling from oracle trajectories (Step 1), forced sampling to collect data (Step 2), and the LCB test (Step 3). In what follows, we further comment on the subtleties of each step.

### v3

@@
Satisficing Regret Minimization in Bandits: Constant Rate and Light-Tailed Distribution
@@
Motivated by the concept of satisficing in decision-making, we consider the problem of satisficing regret minimization in bandit optimization. In this setting, the learner aims at selecting satisficing arms (arms with mean reward exceeding a certain threshold value) as frequently as possible. The performance is measured by satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. We propose SELECT, a general algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence bound Testing, that attains constant expected satisficing regret for a wide variety of bandit optimization problems in the realizable case (i.e., a satisficing arm exists). Specifically, given a class of bandit optimization problems and a corresponding learning oracle with sub-linear expected (standard) regret upper bound, SELECT iteratively makes use of the oracle to identify a potential satisficing arm with low regret. Then, it collects data samples from this arm, and continuously compares the LCB of the identified arm’s mean reward against the threshold value to determine if it is a satisficing arm. As a complement, SELECT also enjoys the same (standard) regret guarantee as the oracle in the non-realizable case. To further ensure stability of the algorithm, we introduce SELECT-LITE that achieves a light-tailed satisficing regret distribution plus a constant expected satisficing regret in the realizable case and a sub-linear expected (standard) regret in the non-realizable case. Notably, SELECT-LITE can operate on learning oracles with heavy-tailed (standard) regret distribution. More importantly, our results reveal the surprising compatibility between constant expected satisficing regret and light-tailed satisficing regret distribution, which is in sharp contrast to the case of (standard) regret. Finally, we conduct numerical experiments to validate the performance of SELECT and SELECT-LITE on both synthetic datasets and a real-world dynamic pricing case study.
@@
online learning, satisficing, regret, tail risk control
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic framework for decision-making under uncertainty. In MAB, a learner is given a set of arms with initially unknown mean rewards. She repeatedly picks an arm to collect its noisy reward and her goal is to maximize the cumulative rewards. The performance is usually measured by the notion of regret, which is the difference between the maximum possible total mean rewards and the total mean rewards collected by the learner. With its elegant form, MAB has found numerous applications in recommendation systems, ads display, and beyond, and many (near-)optimal algorithms have been developed for various classes bandit optimization problems (Lattimore and Szepesvári 2020).
@@
The concept of satisficing finds its place in many real world decision-making under uncertainty settings (Caplin et al. 2011) as shown below.
@@
  • Inventory Management: Studies show that product stockouts can result in customer disengagement and impair a firm’s profitability (Fitzsimons 2000, Jing and Lewis 2011). As such, retailers usually replenish their products periodically to improve service level (i.e., probability of no stockout for any product). Notably, due to various constraints (e.g., capacity and inventory cost), instead of blindly maximizing service level, retailers typically set a target service level, e.g., the ideal service level in retailing is $90\%\sim 95\%$ (Levy et al. 2023). Consequently, in face of the unknown demand distributions, one can cast the problem into a MAB setting with a different objective. In this setting, a retailer starts without complete knowledge of the demand distributions and proceeds by trying various stock level combinations (i.e., arms) of the products over time. Here, the objective is to hit the target service level as frequently as possible rather than maximizing the service level every time.
  • Product Design: In fashion product design, designers usually need to design a sequence of products for a group of customers with initially unknown preferences. Instead of demanding for the best design every time, designers only need to pick a design that can hopefully meet the expectation of a certain portion of the customers for most of the time (Mabey 2022, Whitenton 2014).
@@
The above examples give rise to the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2023) where the objective is to hit a satisficing arm (i.e., an arm whose mean reward exceeds a certain threshold value) as frequently as possible. The corresponding performance measure is satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. In particular, Michel et al. (2023) develops SAT-UCB, an algorithm for finite-armed satisficing bandits: In the realizable case (i.e., a satisficing arm exists), the algorithm attains a constant expected satisficing regret, i.e., independent of the length of the time horizon; Otherwise, in the non-realizable case (i.e., the threshold value is out of reach), it preserves the optimal logarithmic expected (standard) regret.
@@
Despite these, existing algorithms usually impose that there is a clear separation between the threshold value and the largest mean reward among all non-satisficing arms, which is referred to as the satisficing gap. Moreover, the satisficing regret bounds usually scale inverse proportionally to the satisficing gap. On the other hand, it is evident that in many bandit optimization problem classes (e.g., Lipschitz bandits and concave bandits), the satisficing gap can simply be 0, making the bounds vacuous. Therefore, it is not immediately clear if SAT-UCB (Michel et al. 2023) or other prior results can go beyond finite-armed bandits and if constant satisficing regret is possible for general classes of bandit optimization problems.
1.1 Main Contributions
@@
In this work, we propose SELECT, a novel algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence Bound Testing. More specifically:
@@
  • We describe the design of SELECT in Section 3. For a given bandit optimization problem class and an associated learning oracle with sub-linear (but not necessarily optimal) expected standard regret guarantee, SELECT runs in rounds with the help of the oracle. At the beginning of a round, SELECT first applies the oracle for a number of time steps and randomly samples an arm from its trajectory as a candidate satisficing arm. Then, it conducts forced sampling by pulling this arm for a pre-specified number of times. Finally, it continues to pull this arm and starts to monitor the resulted lower confidence bound (LCB) of its mean reward. SELECT terminates the current round and starts the next once the LCB falls below the threshold value;
@@
  • In Section 4, we establish that in the realizable case, i.e., when a satisficing arm exists, SELECT is able to achieve a constant expected satisficing regret. Notably, the expected satisficing regret bound has no dependence on the satisficing gap (see the forthcoming Remark 3 for a detailed discussion). Instead, it scales inverse proportionally to the exceeding gap, which measures the difference between the optimal reward and the threshold value, and is positive in general; Otherwise (i.e., when no satisficing arm exists), SELECT enjoys the same expected standard regret bound as the learning oracle.
@@
  • While SELECT attains constant satisficing regret in expectation, its performance in individual realizations may not be stable, i.e., the probability of incurring a large satisficing regret is not sufficiently small. To address this, in Section 6, we consider the tail risk control in satisficing regret minimization. To confirm the necessity of new algorithms for light-tailed satisficing regret, we first prove that the satisficing regret distribution of SELECT can be heavy-tailed even if we use a learning oracle with light-tailed standard regret distribution. Then, we introduce SELECT-LITE, a modified version of SELECT to obtain a light-tailed satisficing regret distribution. Specifically, with an arbitrary learning oracle with sub-linear expected standard regret bound, SELECT-LITE enjoys the following desirable properties:
    • In the realizable case, SELECT-LITE attains a constant expected satsficing regret bound, and its satisficing regret distribution is light-tailed; and
    • In the non-realizable case, SELECT-LITE still maintains a sub-linear in $T$ standard regret bound.
    Notably, our results do not rely on any assumptions on the tail distribution of the learning oracle’s standard regret. In fact, SELECT-LITE is able to make use of a learning oracle with heavy-tailed standard regret to obtain a light-tailed satisficing regret distribution in the realizable case. Moreover, in Simchi-Levi et al. (2024), it is shown that any algorithm with a light-tailed standard regret distribution has to suffer a $\Omega(\text{poly}(T))$ instance-dependent expected standard regret, rather than a near-optimal $O(\log(T))$ instance-dependent expected standard regret. Because the constant expected satisficing regret bounds are typically instance-dependent, it becomes unclear if a light-tailed satisficing regret distribution will inevitably cause any dependence on $T$ in the expected satisficing regret. Nevertheless, our results show that achieving a light-tailed satisficing regret only worsens the expected satisficing regret by a constant. This thus reveals the compatibility between constant expected satisficing regret and light-tailed satisficing regret distribution. It also indicates a fundamental difference between the notions of standard regret and satisficing regret.
  • In Section 7, we conduct numerical experiments to demonstrate the performance of SELECT on synthetic instances of finite-armed bandits, concave bandits, and Lipschitz bandits. We use SAT-UCB, SAT-UCB+ (a heuristic with no regret guarantee proposed by Michel et al. 2023), and the respective learning oracle as the benchmarks. Our results reveal that in the realizable case, SELECT does achieve constant expected satisficing regret. Moreover, both its expected satisficing regrets (in the realizable cases) and standard regrets (in the non-realizable cases) either significantly outperform the benchmarks’ or are close to the best-performing ones’. We also test the performance of SELECT-LITE and compare it with SELECT. We show that SELECT-LITE has better performance in terms of tail distribution control compared to SELECT while still maintaining a constant expected satisficing regret comparable to SELECT. This indicates that SELECT-LITE can provide better protection against tail risks without incurring a significantly larger satisficing regret in expectation.
  • To test the performance of our algorithm in real-world instances, in Section 8, we also conduct numerical experiments using a case study of dynamic pricing calibrated from a real-world dataset. Using a real-world avocado sales dataset Kiggins (2018), we show that SELECT attains a constant expected satisficing regret in the realizable case. Furthermore, our results reveal that SELECT outperforms all other benchmarks both in the realizable and non-realizable cases.
  • A preliminary version of this paper (Feng et al. 2025) has appeared in the Proceedings of the 13th International Conference on Learning Representations (ICLR 2025). The current paper provides additional contributions in three directions.
    • First, we go beyond minimizing satisficing regret only in expectation and consider the tail risk control in satisficing regret minimization. Specifically, we adapt SELECT and establish SELECT-LITE, which is able to attain a light-tailed satisficing regret distribution (Theorem 5) while still maintaining a constant expected satisficing regret in the realizable case (Theorem 6) for a wide range of bandit problems;
    • Next, in Section 7.4, we test the empirical performance of SELECT-LITE using an instance of finite-armed bandit. The results show that SELECT-LITE is able to maintain a smaller tail probability in satisficing regret compared to SELECT while the expected satisficing regret of SELECT-LITE is comparable to that of SELECT;
    • Finally, in Section 8, we test our algorithm using a case study of dynamic pricing calibrated from a real-world avocado sales dataset. We observe that under this instance SELECT is able to significantly outperform all other benchmarks, demonstrating a strong empirical performance of our algorithm in practical bandit problems.
@@ 1.2 Related Works
Another line of research relevant to our work is tail risk control in bandits. The general goal of this line of research is to design bandit algorithms that achieve light-tailed regret distributions besides minimizing expected regret. Some early works studying the tail distribution of bandit algorithms include Audibert et al. (2009) and Salomon and Audibert (2011), who show that for UCB algorithms, the probability of regret greater than $c(\log(T))^{p}$ (where $p>1$ is fixed) decays only polynomially in $T$. Fan and Glynn (2024) are among the first to establish bandit algorithms with a light-tailed (standard) regret distribution. They focus on the finite-armed bandit setting and show that for informationally-optimized bandit algorithms such as the UCB algorithm and Thompson sampling, the probability of incurring a linear in $T$ regret is at least $\Omega(1/T)$. Some follow-up works include Simchi-Levi et al. (2023) and Simchi-Levi et al. (2024). They show that any finite-armed bandit algorithm with a $O(\log(T))$ instance-dependent expected regret bound has a heavy-tailed regret distribution. They also establish an optimal tradeoff between tail risk, worst-case expected regret bound and instance-dependent expected regret bound. In our paper, we consider tail risk control for satisficing regret minimization. We show that unlike tail risk control in standard regret minimization, where $O(\log(T))$ instance-dependent expected regret bound cannot coexist with light-tailed regret distribution, in satisficing regret minimization there exists algorithm that attains a constant expected satisficing regret and a light-tailed satisficing regret distribution.
@@ 2 Problem Formulation
In this section, we present the setup of our problem. Consider a class of bandit optimization problems $(\mathcal{A},\mathcal{R})$, where $\mathcal{A}$ is the arm set and $\mathcal{R}\subseteq\{r:\mathcal{A}\to[0,1]\}$ is the class of admissible reward functions. The learner is given a set of feasible arms $\mathcal{A}$ and a class of admissible reward functions $\mathcal{R}$, pulling arm $A\in\mathcal{A}$ in a time step brings a random reward with mean $r(A)$. The learner knows the underlying reward function $r(\cdot)$ belongs to the function class $\mathcal{R}$, but the exact $r(\cdot)$ is initially unknown to the learner.
In each time step $t=1,2,\ldots,T$, the learner pulls an arm $A_{t}\in\mathcal{A}$, and then observes a noisy reward $Y_{t}=r(A_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is a conditionally 1-subgaussian random variable, i.e., $\mathbb{E}[\epsilon_{t}|A_{1},\epsilon_{1}\ldots,A_{t-1},\epsilon_{t-1},A_{t}]=0$ and $\mathbb{E}[e^{\lambda\epsilon_{t}}|A_{1},\epsilon_{1}\ldots,A_{t-1},\epsilon_{t-1},A_{t}]\leq e^{\lambda^{2}/2}$ holds for all $\lambda\in\mathbb{R}$.
@@
$\texttt{Regret}_{S}=\sum_{t=1}^{T}\max\{S-r(A_{t}),0\}.$
Because it is possible that no arm achieves the satisficing level, we also follow the conventional bandit literature to define the (standard) regret, which measures the expected cumulative difference of between the chosen arm’s mean reward and the optimal mean reward. Specifically, let $A^{*}=\arg\max_{A\in\mathcal{A}}r(A),$ the standard regret is defined as
$\texttt{Regret}=T\cdot r(A^{*})-\sum_{t=1}^{T}r(A_{t}).$
We say that the setting is realizable if $r(A^{*})\geq S$; Otherwise, it is non-realizable. We recall that an arm $A$ is satisficing if $r(A)\geq S$; Otherwise, it is non-satisficing.
Remark 2.
In the definition of satisficing regret, we consider the cumulative deficit of mean reward compared to the satisficing threshold because ideally we would like the arm pull in every time step to be satisficing, thus we penalize every arm pull of non-satisficing arms. Even though in some time steps the decision maker may pull an arm with mean reward over the satisficing threshold, the excessive mean reward cannot compensate for the satisficing regret incurred in other time steps. With this definition of satisficing regret, the problem cannot be trivially solved by running a traditional bandit algorithm. Although any bandit algorithm will gradually converge to the optimal arm and achieve a satisficing total reward with long enough time horizon, as time horizon $T$ increase the number of arm pulls of non-satisficing arm will also increase, incurring a satisficing regret scaling in $T$. In the forthcoming Section 7 and Section 8, we also numerically show that directly running bandit algorithms can incur substantial satisficing regret.
Existing works (e.g., Michel et al. 2023) are mostly focused on satisficing regret minimization in finite-armed bandit settings, and usually derive expected satisficing regret bound that scales inversely proportional to the satisficing gap
$\Delta_{S}=\min\{S-r(A):r(A)<S\}.$
@@ 3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
In this section, we present our algorithmic template SELECT for general classes of bandit optimization problems. When the arm set $\mathcal{A}$ is not too large, one can collect enough data samples for every arm to estimate its mean reward and gradually abandon any non-satisficing arms (see, e.g., Garivier et al. 2019, Michel et al. 2023). In general, however, the arm set can be large and may even contain infinitely many arms (e.g., continuous arm set). Therefore, it becomes impossible to identify all non-satisficing arms. We thus take an alternative approach: The algorithm repeatedly locates a potentially satisficing arm with low (satisficing) regret. Then, it tests if this candidate arm is truly a satisficing arm or not. If not, it kicks off the next round of search.
To ensure fast detection of candidate arms with low regret for a given bandit optimization problem class $(\mathcal{A},\mathcal{R})$, SELECT makes use of a bandit algorithm for (standard) regret minimization on $(\mathcal{A},\mathcal{R})$. To this end, we assume access to a blackbox learning oracle that achieves sub-linear standard regret for the bandit optimization problem class $(\mathcal{A},\mathcal{R})$.
@@
For problem class $(\mathcal{A},\mathcal{R})$, there exists a sequence of learning algorithms ALG such that for any reward function $r\in\mathcal{R}$ and any time horizon $t\geq 2$, the expected regret of $\texttt{ALG}(t)$ is bounded by
$t\cdot r(A^{*})-\mathbb{E}\left[\sum_{s=1}^{t}r(A_{s})\right]\leq C_{1}t^{\alpha}\log(t)^{\beta}.$
@@
We remark that Condition 1 only asks for sub-linearity in standard regret upper bounds rather than optimality. Therefore, it is an extremely mild condition and can be easily satisfied by a wide range of bandit optimization problem classes and the corresponding bandit learning oracles, e.g., finite-armed bandits, linear bandits, concave bandits, Lipschitz bandits, etc. We will demonstrate this in the forthcoming Section 5 and Section I of the Appendix.
With this, for a problem class $(\mathcal{A},\mathcal{R})$ and its sub-linear regret learning algorithm ALG, SELECT runs in rounds. In round $i$, it takes the following steps (see Fig. 1 for a illustration):
@@
Step 1. Identify a Potentially Satisficing Arm: Let $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$, SELECT follows $\texttt{ALG}(t_{i})$ for the first $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil=\lceil 2^{i/\alpha}\rceil$ steps of round $i$ and records its arm selections. Then, it samples an arm $\hat{A}_{i}$ uniformly random from the trajectory of $\texttt{ALG}(t_{i})$ in this round. By virtue of ALG, we find an arm whose mean reward is at most $\tilde{O}(t^{\alpha-1}_{i})=\tilde{O}(\gamma_{i})$ below $r(A^{*})$ in expectation, and only $\tilde{O}(t^{\alpha}_{i})$ satisficing regret is incurred in this step in the realizable case [^1] We adopt the asymptotic notations $O(\cdot),\Omega(\cdot),$ and $\Theta(\cdot)$ as defined in (Cormen et al. 2009). When logarithmic factors are omitted, we write them as $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),$ and $\tilde{\Theta}(\cdot).$. We note that in the realizable case, as $i$ increases, $\tilde{O}(\gamma_{i})$ will gradually become smaller than $\Delta_{S}^{*}$, meaning that the sampled arm $\hat{A}_{i}$ is more likely to be a satisficing arm and enjoys no satisficing regret;
Step 2. Forced Sampling on the Identified Arm: To validate if $\hat{A}_{i}$ is a satisficing arm or not, we need to collect enough data from pulling $\hat{A}_{i}$ before performing any statistical test (see the forthcoming Remark 3 for the reasons). However, when the arm set $\mathcal{A}$ is large, $\texttt{ALG}(t_{i})$ might only have pulled $\hat{A}_{i}$ for limited number of times. To circumvent this, we perform a forced sampling on $\hat{A}_{i}$ by pulling it for $T_{i}=\lceil\gamma^{-2}_{i}\rceil$ times;
Step 3. Lower Confidence Bound Tests: In this step, SELECT kicks off the LCB test as more data is collected from pulling $\hat{A}_{i}$. We use $k$ (initialized to 0) to denote the number of additional pieces of noisy reward data generated by pulling $\hat{A}_{i}$ in the current time step and $\hat{r}^{\text{tot}}_{i}$ to denote the running total reward collected from Step 2 and the current step so far. At the same time, it keeps comparing the LCB of $\hat{A}_{i}$’s mean reward, i.e., $\hat{r}^{\text{tot}}_{i}/(T_{i}+k)-\sqrt{4\log(T_{i}+k)/(T_{i}+k)}$, against the threshold value $S$ after acquiring every piece of new data. SELECT terminates this round and enters the next whenever the LCB is less than $S$.
@@
Remark 3 (Key Novelties of SELECT).
Our algorithm brings together several key ingredients, including sampling from oracle trajectories (Step 1), forced sampling to collect data (Step 2), and the LCB test (Step 3). In what follows, we further comment on the subtleties of each step.